OpenStudy (anonymous):

4 1/8 - 2 7/8 + 2 1/4 =

OpenStudy (anonymous):

find common denominators first, then subtract/add two numbers at a time

OpenStudy (texaschic101):

4 1/8 - 2 7/8 + 2 1/4 (turn to improper fractions) 33/8 - 23/8 + 9/4 10/8 + 9/4 = 5/4 + 9/4 = 14/4 = 3 1/2 or 33/8 - 23/8 + 9/4 10/8 + 9/4 = 10/8 + 18/8 = 28/8 = 3 1/2